OpenStudy (anonymous):

Let f(x) = x2 – 16 and g(x) = x + 4. Find f/g and its domain. help please? :/

jimthompson5910 (jim_thompson5910):

can you factor x^2 - 16 ?

OpenStudy (anonymous):

Not sure how to do it on my ipod calculator.

jimthompson5910 (jim_thompson5910):

are you familiar with the difference of squares rule?

jimthompson5910 (jim_thompson5910):

if not, then a^2 - b^2 = (a + b)(a - b)

jimthompson5910 (jim_thompson5910):

so how can you use this to factor x^2 - 16 ?

OpenStudy (anonymous):

x + 4; all real numbers except x ≠ 4 x + 4; all real numbers except x ≠ –4 x – 4; all real numbers except x ≠ 4 x – 4; all real numbers except x ≠ –4 these are the choices.

jimthompson5910 (jim_thompson5910):

ok, but how can you use a^2 - b^2 = (a + b)(a - b) to factor x^2 - 16 ?

OpenStudy (skullpatrol):

Do you know why you can never divide by 0?